Suppose our underlying is a stock XYZ. Today (t=0), XYZ is priced at $1,070. The storage and insurance cost is $13, paid in advance. The forward contract uses XYZ as the underlying, which will expire in one year from today. The interest rate is 0.044. The forward price at today (t=0) is $1,441. What is the arbitrage profit that you can make today based on cost-of-carry model?
Answer #1

Price of the forward should be=(1070+13)*(1+0.044)=$1,130.65

As forward is overpriced, sell the forward buy the spot and borrow money

Net cash flow at t=0: Borrow 1083 Buy spot and pay for storage cost cash flow=1083-1070-13=0
Net cash flow at t=1: 1441-1083*1.044=310.348 Deliver the spot to forward buyer

Net profit=310.348